# Renormalization Group is the Principle behind the Holographic Entropy Cone

### v1

@@ 1 Introduction
A major theme in the AdS/CFT correspondence Maldacena:1997re; Witten:1998qj is that the gravitational side of the duality is an information theoretic rendering of the field theory side, geometrically organized VanRaamsdonk:2010pw. Key tenets of this information theoretic rendering are:
@@
    The Ryu-Takayanagi proposal Ryu:2006bv; rt2: von Neumann entropies of spatial subregions are manifested as areas of extremal surfaces plus $\mathcal{O}(1/N)$ corrections. We refer to the requisite surfaces as RT surfaces after Ryu and Takayanagi.
@@
    Subregion duality Czech:2012bh; Dong:2016eik: the physics in a boundary spatial subsystem $A$ solely controls the physics in a bulk subsystem called the entanglement wedge ${\rm EW}(A)$, which is demarcated by the Ryu-Takayanagi extremal surface.
@@
    Low energy bulk physics as an erasure correcting code Almheiri:2014lwa: when entanglement wedges of different subsystems overlap, different parties can share control over the same code subspace. That is, overlapping ${\rm EW}(X_{i})$ indicate that parties in boundary regions $X_{i}$ share access to the underlying bulk physics, which in turn is protected against erasures of $\overline{X_{i}}$ (the complement of $X_{i}$).
@@
    Holographic Renormalization Group (RG) Balasubramanian:1999jd: the bulk geometry is organized into layers that correspond to CFT scales, with the near-boundary regions encoding the CFT ultraviolet (UV) physics and the central region encoding the infrared (IR).
In this list, points 1.-3. are directly controlled by quantum entanglement in the CFT—including the erasure correcting property, whose geometric contours are decided by the RT surfaces. Point 4. may seem more distant, but in fact the Renormalization Group is closely interlaced with quantum entanglement. This is evident in the entanglement-based proofs of the $c$- and $F$- and $a$-theorems Casini:2004bw; Myers:2010xs; Casini:2012ei; Casini:2016udt; Casini:2017vbe, measures of UV-IR entanglement Balasubramanian:2011wt and—perhaps most visually striking—in the Multi-scale Entanglement Renormalization Ansatz Vidal:2006sxo; Vidal:2008zz.
@@
While entanglement nesting naturally combines the perspectives of quantum entanglement, renormalization group and geometry, the manner in which it does so is somewhat coarse. Because enlarging $X_{j}$ to obtain $X_{i}\supset X_{j}$ entails adding UV degrees of freedom, ${\rm EW}(X_{i})$ is guaranteed to be larger than ${\rm EW}(X_{j})$ in its boundary extent but may or may not reach deeper in the bulk. It would be desirable to formulate analogues or extensions of entanglement wedge nesting, which apply to regions of equal boundary support (independent of UV entanglement) and specifically probe the bulk depth of entanglement wedges (depend on exclusively IR entanglement). Ideally, such statements would also tell us more detailed information about access structures of holographic erasure correcting codes—which combinations of parties can / cannot decode the same logical information. In this paper, we explain that such statements are already known and well studied, albeit in a somewhat disguised form. They comprise the holographic entropy cone Bao:2015bfa.
The holographic entropy cone collects linear inequalities, which are obeyed by areas of RT surfaces by virtue of their minimality. [^1] This assumes time reversal symmetry. In this paper we sidestep the issue of whether the same inequalities hold when the bulk is not time reflection-symmetric and entanglement entropies are computed by the HRT proposal hrt; maximin. Progress on this question includes Czech:2019lps; Grado-White:2024gtx; Grado-White:2025jci; Grimaldi:2025jad. The cone terminology reflects the fact that each inequality is saturated on a hyperplane in entropy space (the vector space of hypothetical values of subsystem entropies) and, therefore, the set of holographically admissible entropy assignments forms a cone. The inequalities that bound the holographic entropy cone are famously intricate and difficult to parse HernandezCuenca:2019wgh; Hernandez-Cuenca:2023iqh; Czech:2023xed; see ineq. (57) for an example. It has been a long standing challenge to interpret them in physical or operational terms; see Czech:2021rxe; Czech:2025jnw for earlier partial results. We venture to propose an interpretation.
We argue that the holographic entropy inequalities—the bounding facets of the holographic entropy cone—enforce and protect holographic RG. In essence, they stipulate that the more infrared regions’ entanglement wedges reach deeper into the bulk than do entanglement wedges of ultraviolet regions, with the saturation of an inequality implying precisely equal depth. When this rule of thumb is further refined using technical tools, one discovers that saturating an inequality eliminates many otherwise reasonable access structures of the quantum erasure correcting code Czech:2025jnw. Thus, the non-negative quantity defined by each inequality is an order parameter for a certain type of erasure correction.
@@
We consider time reflection-symmetric, asymptotically AdS geometries, which are dual to states of a conformal field theory (CFT). At leading order in Newton’s constant $G_{N}$, the entanglement entropy of a boundary region $X$ is given by Ryu:2006bv; rt2:
@@
Proposal (3) implies that the CFT entanglement entropies satisfy infinitely many inequalities Bao:2015bfa of the schematic form
@@ 1.1 Overview of results
We argue that the inequalities serve the following purpose: to enforce and protect the layered organization of the bulk AdS geometry by CFT scale, with infrared (IR) physics residing deeper in the bulk. This is encapsulated by the following statements, which apply to every inequality (4): [^2] The inequality has to be written in a balanced form He:2020xuo. We explain this point further in Section 2.1.
@@
In prior work Czech:2025jnw, Wang and we showed that the saturation of an inequality ${\rm LHS}={\rm RHS}$ is a no-go condition for certain access structures in a holographic erasure correcting code. Applied directly to inequalities (4), those results are distinct and logically independent from claim (7). However, applying Czech:2025jnw to the tautologically extended inequality ${\rm LHS}+S(Y_{j})\geq{\rm RHS}+S(Y_{j})$ reveals a wealth of further constraints on holographic erasure correction, which apply when (4) is saturated. These constraints provide an improvement on (7), which in many cases functions at the level of individual $E(Y_{j})$ and not their union:
@@ 2.1 Inclusion
Statement (5) can be justified in two ways. The simpler way, which we present here, relies on a conjectured property of all inequalities (4), which was recently proposed in Grimaldi:2025jad. Another argument exploits an auxiliary object called a contraction map, which is used in the conventional technique of proving holographic inequalities Bao:2015bfa. We explain contraction proofs in Appendix A and give a contraction-based proof of (5) in Appendix B.
It is conceivable that some hitherto unknown entropy inequality might be holographically valid and disprove the conjecture of Grimaldi:2025jad and have no proof by the contraction method, though Reference Bao:2025sjn argued that no such inequality exists. We return to this theoretical possibility in the Discussion.
@@
The conjecture in Grimaldi:2025jad stipulates that every holographic inequality passes a test, which we presently explain. We use the monogamy of mutual information Hayden:2011ag
@@
The test presumes that the inequality is written in a balanced form—one where every region appears the same number of times on the LHS and on the RHS. [^3] Working with global pure states and assuming Haag duality holds (see e.g. Casini:2019kex for scenarios where it does not), it is possible to rewrite a balanced inequality in an unbalanced form by exploiting $S(X)=S(\overline{X})$. Such rewritings can sometimes be useful Czech:2023xed, but not for the purposes of this paper. Example (10) manifestly satisfies the balanced requirement. Reference He:2020xuo showed that every holographic entropy inequality can be written in a balanced form.
@@
One retort is that they are necessary to render the difference $({\rm LHS}-{\rm RHS})$ UV-finite in the CFT. Indeed, UV-finiteness—known under the monicker ‘superbalance’—is a proven property of all holographic entropy inequalities He:2020xuo. But superbalance entails a further corollary: the inequality must pass the majorization test of Grimaldi:2025jad even if we rewrite it using a different choice of purifier. We explain this presently.
@@
This rewriting treats $C$ as the purifier—that is, we exploit $S(X)=S(\overline{X})$ to rewrite every term in the way that excludes $C$. When we change purifiers in a superbalanced inequality, the rewritten inequality still passes the test of Grimaldi:2025jad and, consequently, respects claims (5) and (6). This is not true for inequalities, which are not superbalanced.
@@ 2.2 Saturation
We work with inequalities (4) that can be proven by the contraction method Bao:2015bfa; if a valid inequality not provable by contraction exists, our results do not apply to it. We review proofs by contraction in Appendix A, but the argument can be stated without reference to the mechanics of such proofs. The following text is self-contained in that a working knowledge of contraction maps is not necessary.
@@
The only caveat in this argument is if two surfaces $\Sigma^{\rm can}_{j}$ and $\Sigma^{\prime}_{j}$ simultaneously achieve the minimum in (3). We would then be at a phase transition—a non-generic scenario, in which proposal (3) is subject to enhanced corrections Dong:2020iod; Akers:2020pmf. A common prescription for defining entanglement wedges at a phase transition is to select the smallest possible wedge $E(Y_{j})\subseteq E^{\rm can}(Y_{j})$, which again confirms (7).
@@ 2.3 Examples
Thus, strong subadditivity evades the result of Czech:2025jnw and it is not superbalanced. Consequently, we do not get to apply (7) for every choice of purifier in (15). However, strong subadditivity is symmetric under the exchange $B\leftrightarrow O$ (where the state on $ABCO$ is pure) so we have:
@@
Conclusion (16)—or rather its contrapositive—was previously established in May:2021raz. That analysis was more robust because it did not assume time reversal symmetry and applied to full entanglement wedges, not merely their restrictions to a specific bulk slice.
@@
Figure 2 illustrates results (22) and (24) using consecutively labeled intervals on the boundary of a static BTZ geometry Banados:1992wn. The figure manifests the main qualitative point of the paper: When an inequality is saturated, the more ‘infrared’ entanglement wedges fail to reach an incremental depth in the bulk.
@@ 3 A refinement
In Reference Czech:2025jnw, Wang and we explained that contraction proofs of holographic entropy inequalities strongly constrain what happens when the said inequalities are saturated. When those results are applied directly to inequality (4), they are distinct and logically independent from claims (5-7). However, a trick unifies the two approaches. If, instead of applying Czech:2025jnw to inequality ${\rm LHS}\geq{\rm RHS}$, one applies it to
@@
In this section—and only here—we assume knowledge of proofs by contraction and of the results derived in Reference Czech:2025jnw. To make the paper self-contained, we review proofs by contraction in Appendix A and the findings of Czech:2025jnw in Appendix C. In order to establish common ground, we list below the necessary facts about contraction maps and implications of saturation—without justification. All facts highlighted below are justified in the Appendices.
@@ 3.1 Key facts
Saturation implies that nesting-violating $W(x)$ are empty Czech:2025jnw
@@ 3.2 Example: Subadditivity
The $x=11$ row is greyed out because it refers to the automatically empty region $W(11)=E(A)\cap E(B)$. We observed that such bit strings—characterized in equations (67)—are removed from the domain of the contraction Avis:2021xnz; lhsconflicts. In what follows we do not write out such bit strings in contraction tables except when emphasizing that they are unphysical.
@@ 3.4 The refinement
By applying conclusions (35-36) to bit strings $x$ other than $x=00\ldots 00$, it is sometimes possible to significantly strengthen result (7). In this subsection we exhibit this refinement as it arises in the infinite families of holographic entropy inequalities Czech:2023xed.
@@ 3.4.1 Dihedral inequalities
The dihedral inequalities Bao:2015bfa take the form:
@@
In particular, the highlighted values of $f(x)_{j}$ persist for all choices of bits $x_{i}$, which are left undetermined (appear as ‘any’) in Table 5. These properties were proven for all values of $m$ in Bao:2015bfa. It is amusing to confirm them with the explicit contraction map for $m=5$, which is found in Appendix B of Bao:2015bfa.
Apply (35) to bit strings $x$ in the second row of Table 5. We find that if (46) is saturated then:
$E(A_{1}^{+}O)\subseteq\overline{E(A_{1}^{+})}\cap\overline{E(A_{1}^{-}O)}\cap\overline{E(A_{2}^{-}O)}\cap\ldots$ (47)
@@
$E(A_{1}^{+}O)\subseteq\overline{E(A_{1}^{+})}\cap\overline{E(A_{1}^{-}O)}\cap\overline{E(A_{2}^{-}O)}=E(A_{1}^{+})\cup E(A_{1}^{-}O)\cup E(A_{2}^{-}O)$ (48)
@@ 3.4.2 Toric inequalities
In Czech:2022fzb; Czech:2023xed; Czech:2024rco, the following generalization of the dihedral inequalities was proposed and proven:
@@
As shown in Czech:2023xed, contraction proofs of inequality (52) exhibit the following pattern:
@@
In References Czech:2023xed; Czech:2025tds, we introduced a graphical representation of bit strings $x$, which involves nonintersecting loops living on an $n\times m$-square tessellation of a torus. The rows highlighted in Table 6—and the refined statement (55)—arise from those bit strings, which are represented by minimal loops.

### v2

@@ 1 Introduction
A major theme in the AdS/CFT correspondence Maldacena (1998); Witten (1998) is that the gravitational side of the duality is an information theoretic rendering of the field theory side, geometrically organized Van Raamsdonk (2010). Key tenets of this information theoretic rendering are:
@@
    The Ryu-Takayanagi proposal Ryu and Takayanagi (2006b, a): von Neumann entropies of spatial subregions are manifested as areas of extremal surfaces plus $\mathcal{O}(1/N)$ corrections. We refer to the requisite surfaces as RT surfaces after Ryu and Takayanagi.
@@
    Subregion duality Czech et al. (2012); Dong et al. (2016): the physics in a boundary spatial subsystem AA solely controls the physics in a bulk subsystem called the entanglement wedge ${\rm EW}(A)$, which is demarcated by the Ryu-Takayanagi extremal surface.
@@
    Low energy bulk physics as an erasure correcting code Almheiri et al. (2015): when entanglement wedges of different subsystems overlap, different parties can share control over the same code subspace. That is, overlapping ${\rm EW}(X_{i})$ indicate that parties in boundary regions $X_{i}$ share access to the underlying bulk physics, which in turn is protected against erasures of $\overline{X_{i}}$ (the complement of $X_{i}$).
@@
    Holographic Renormalization Group (RG) Balasubramanian and Kraus (1999): the bulk geometry is organized into layers that correspond to CFT scales, with the near-boundary regions encoding the CFT ultraviolet (UV) physics and the central region encoding the infrared (IR).
In this list, points 1.-3. are directly controlled by quantum entanglement in the CFT—including the erasure correcting property, whose geometric contours are decided by the RT surfaces. Point 4. may seem more distant, but in fact the Renormalization Group is closely interlaced with quantum entanglement. This is evident in the entanglement-based proofs of the $c$- and $F$- and $a$-theorems Casini and Huerta (2004); Myers and Sinha (2010); Casini and Huerta (2012); Casini et al. (2017b, a), measures of UV-IR entanglement Balasubramanian et al. (2012) and—perhaps most visually striking—in the Multi-scale Entanglement Renormalization Ansatz Vidal (2007, 2008).
@@
While entanglement nesting naturally combines the perspectives of quantum entanglement, renormalization group and geometry, the manner in which it does so is somewhat coarse. Because enlarging $X_{j}$ to obtain $X_{i}\supset X_{j}$ entails adding UV degrees of freedom, ${\rm EW}(X_{i})$ is guaranteed to be larger than ${\rm EW}(X_{j})$ in its boundary extent but may or may not reach deeper in the bulk. It would be desirable to formulate analogues or extensions of entanglement wedge nesting, which apply to regions of equal boundary support (independent of UV entanglement) and specifically probe the bulk depth of entanglement wedges (depend on exclusively IR entanglement). Ideally, such statements would also tell us more detailed information about access structures of holographic erasure correcting codes—which combinations of parties can / cannot decode the same logical information. In this paper, we explain that such statements are already known and well studied, albeit in a somewhat disguised form. They comprise the holographic entropy cone Bao et al. (2015).
The holographic entropy cone collects linear inequalities, which are obeyed by areas of RT surfaces by virtue of their minimality. [^1] This assumes time reversal symmetry. In this paper we sidestep the issue of whether the same inequalities hold when the bulk is not time reflection-symmetric and entanglement entropies are computed by the HRT proposal Hubeny et al. (2007); Wall (2014). Progress on this question includes Czech and Dong (2019); Grado-White et al. (2025b, a); Grimaldi et al. (2025). The cone terminology reflects the fact that each inequality is saturated on a hyperplane in entropy space (the vector space of hypothetical values of subsystem entropies) and, therefore, the set of holographically admissible entropy assignments forms a cone. The inequalities that bound the holographic entropy cone are famously intricate and difficult to parse Hernández Cuenca (2019); Hernández-Cuenca et al. (2024); Czech et al. (2024b); see ineq. (57) for an example. It has been a long standing challenge to interpret them in physical or operational terms; see Czech and Shuai (2022); Czech et al. (2025) for earlier partial results. We venture to propose an interpretation.
We argue that the holographic entropy inequalities—the bounding facets of the holographic entropy cone—enforce and protect holographic RG. In essence, they stipulate that the more infrared regions’ entanglement wedges reach deeper into the bulk than do entanglement wedges of ultraviolet regions, with the saturation of an inequality implying precisely equal depth. When this rule of thumb is further refined using technical tools, one discovers that saturating an inequality eliminates many otherwise reasonable access structures of the quantum erasure correcting code Czech et al. (2025). Thus, the non-negative quantity defined by each inequality is an order parameter for a certain type of erasure correction.
@@
We consider time reflection-symmetric, asymptotically AdS geometries, which are dual to states of a conformal field theory (CFT). At leading order in Newton’s constant $G_{N}$, the entanglement entropy of a boundary region $X$ is given by Ryu and Takayanagi (2006b, a):
@@
Proposal (3) implies that the CFT entanglement entropies satisfy infinitely many inequalities Bao et al. (2015) of the schematic form
@@ 1.1 Overview of results
We argue that the inequalities serve the following purpose: to enforce and protect the layered organization of the bulk AdS geometry by CFT scale, with infrared (IR) physics residing deeper in the bulk. This is encapsulated by the following statements, which apply to every inequality (4): [^2] The inequality has to be written in a balanced form He et al. (2020). We explain this point further in Section 2.1.
@@
In prior work Czech et al. (2025), Wang and we showed that the saturation of an inequality ${\rm LHS}={\rm RHS}$ is a no-go condition for certain access structures in a holographic erasure correcting code. Applied directly to inequalities (4), those results are distinct and logically independent from claim (7). However, applying Czech et al. (2025) to the tautologically extended inequality ${\rm LHS}+S(Y_{j})\geq{\rm RHS}+S(Y_{j})$ reveals a wealth of further constraints on holographic erasure correction, which apply when (4) is saturated. These constraints provide an improvement on (7), which in many cases functions at the level of individual $E(Y_{j})$ and not their union:
@@ 2.1 Inclusion
Statement (5) can be justified in two ways. The simpler way, which we present here, relies on a conjectured property of all inequalities (4), which was recently proposed in Grimaldi et al. (2025). Another argument exploits an auxiliary object called a contraction map, which is used in the conventional technique of proving holographic inequalities Bao et al. (2015). We explain contraction proofs in Appendix A and give a contraction-based proof of (5) in Appendix B.
It is conceivable that some hitherto unknown entropy inequality might be holographically valid and disprove the conjecture of Grimaldi et al. (2025) and have no proof by the contraction method, though Reference Bao et al. (2025) argued that no such inequality exists. We return to this theoretical possibility in the Discussion.
@@
The conjecture in Grimaldi et al. (2025) stipulates that every holographic inequality passes a test, which we presently explain. We use the monogamy of mutual information Hayden et al. (2013)
@@
The test presumes that the inequality is written in a balanced form—one where every region appears the same number of times on the LHS and on the RHS. [^3] Working with global pure states and assuming Haag duality holds (see e.g. Casini et al. (2020) for scenarios where it does not), it is possible to rewrite a balanced inequality in an unbalanced form by exploiting $S(X)=S(\overline{X})$. Such rewritings can sometimes be useful Czech et al. (2024b), but not for the purposes of this paper. Example (10) manifestly satisfies the balanced requirement. Reference He et al. (2020) showed that every holographic entropy inequality can be written in a balanced form.
@@
One retort is that they are necessary to render the difference $({\rm LHS}-{\rm RHS})$ UV-finite in the CFT. Indeed, UV-finiteness—known under the monicker ‘superbalance’—is a proven property of all holographic entropy inequalities He et al. (2020). But superbalance entails a further corollary: the inequality must pass the majorization test of Grimaldi et al. (2025) even if we rewrite it using a different choice of purifier. We explain this presently.
@@
This rewriting treats $C$ as the purifier—that is, we exploit $S(X)=S(\overline{X})$ to rewrite every term in the way that excludes $C$. When we change purifiers in a superbalanced inequality, the rewritten inequality still passes the test of Grimaldi et al. (2025) and, consequently, respects claims (5) and (6). This is not true for inequalities, which are not superbalanced.
@@ 2.2 Saturation
We work with inequalities (4) that can be proven by the contraction method Bao et al. (2015); if a valid inequality not provable by contraction exists, our results do not apply to it. We review proofs by contraction in Appendix A, but the argument can be stated without reference to the mechanics of such proofs. The following text is self-contained in that a working knowledge of contraction maps is not necessary.
@@
The only caveat in this argument is if two surfaces $\Sigma^{\rm can}_{j}$ and $\Sigma^{\prime}_{j}$ simultaneously achieve the minimum in (3). We would then be at a phase transition—a non-generic scenario, in which proposal (3) is subject to enhanced corrections Dong and Wang (2020); Akers and Penington (2021). A common prescription for defining entanglement wedges at a phase transition is to select the smallest possible wedge $E(Y_{j})\subseteq E^{\rm can}(Y_{j})$, which again confirms (7).
@@ 2.3 Examples
Thus, strong subadditivity evades the result of Czech et al. (2025) and it is not superbalanced. Consequently, we do not get to apply (7) for every choice of purifier in (15). However, strong subadditivity is symmetric under the exchange $B\leftrightarrow O$ (where the state on $ABCO$ is pure) so we have:
@@
Conclusion (16)—or rather its contrapositive—was previously established in May (2021). That analysis was more robust because it did not assume time reversal symmetry and applied to full entanglement wedges, not merely their restrictions to a specific bulk slice.
@@
Figure 2 illustrates results (22) and (24) using consecutively labeled intervals on the boundary of a static BTZ geometry Banados et al. (1992). The figure manifests the main qualitative point of the paper: When an inequality is saturated, the more ‘infrared’ entanglement wedges fail to reach an incremental depth in the bulk.
@@ 3 A refinement
In Reference Czech et al. (2025), Wang and we explained that contraction proofs of holographic entropy inequalities strongly constrain what happens when the said inequalities are saturated. When those results are applied directly to inequality (4), they are distinct and logically independent from claims (5-7). However, a trick unifies the two approaches. If, instead of applying Czech et al. (2025) to inequality ${\rm LHS}\geq{\rm RHS}$, one applies it to
@@
In this section—and only here—we assume knowledge of proofs by contraction and of the results derived in Reference Czech et al. (2025). To make the paper self-contained, we review proofs by contraction in Appendix A and the findings of Czech et al. (2025) in Appendix C. In order to establish common ground, we list below the necessary facts about contraction maps and implications of saturation—without justification. All facts highlighted below are justified in the Appendices.
@@ 3.1 Key facts
Saturation implies that nesting-violating $W(x)$ are empty Czech et al. (2025)
@@ 3.2 Example: Subadditivity
The $x=11$ row is greyed out because it refers to the automatically empty region $W(11)=E(A)\cap E(B)$. We observed that such bit strings—characterized in equations (67)—are removed from the domain of the contraction Avis and Hernández-Cuenca (2023); Li et al. (2022). In what follows we do not write out such bit strings in contraction tables except when emphasizing that they are unphysical.
@@ 3.4 The refinement
By applying conclusions (35-36) to bit strings $x$ other than $x=00\ldots 00$, it is sometimes possible to significantly strengthen result (7). In this subsection we exhibit this refinement as it arises in the infinite families of holographic entropy inequalities Czech et al. (2024b).
@@ 3.4.1 Dihedral inequalities
The dihedral inequalities Bao et al. (2015) take the form:
@@
In particular, the highlighted values of $f(x)_{j}$ persist for all choices of bits $x_{i}$, which are left undetermined (‘any’) in Table 5. These properties were proven for all values of $m$ in Bao et al. (2015). It is amusing to confirm them with the explicit contraction map for $m=5$, which is found in Appendix B of Bao et al. (2015).
Apply (35) to bit strings $x$ in the second row of Table 5. We find that if (46) is saturated then
$E(A_{1}^{+}O)\cap\overline{E(A_{1}^{+})}\cap\overline{E(A_{1}^{-}O)}\cap\overline{E(A_{2}^{-}O)}\cap\ldots=\emptyset$ (47)
@@
$E(A_{1}^{+}O)\subseteq\overline{\overline{E(A_{1}^{+})}\cap\overline{E(A_{1}^{-}O)}\cap\overline{E(A_{2}^{-}O)}\phantom{\Big|}}=E(A_{1}^{+})\cup E(A_{1}^{-}O)\cup E(A_{2}^{-}O)$ (48)
@@ 3.4.2 Toric inequalities
In Czech and Wang (2023); Czech et al. (2024b, a), the following generalization of the dihedral inequalities was proposed and proven:
@@
As shown in Czech et al. (2024b), contraction proofs of inequality (52) exhibit the following pattern:
@@
In References Czech et al. (2024b); Czech and Shuai (2025), we introduced a graphical representation of bit strings $x$, which involves nonintersecting loops living on an $n\times m$-square tessellation of a torus. The rows highlighted in Table 6—and the refined statement (55)—arise from those bit strings, which are represented by minimal loops.